OpenStudy (anonymous):

Find x if y = –1. 6 – 3xy = 12

OpenStudy (jgirl128):

6 - 3x(-1) = 12 simplify 6 - (-3x) = 12 6 + 3x = 12 subtract 6 from both sides 3x = 6 ___ ___ 3 3 x = 2

OpenStudy (anonymous):

not bad , thanks !

OpenStudy (jgirl128):

do you understand?

OpenStudy (jgirl128):

no problem :)

OpenStudy (anonymous):

yeah , i just thought that you wouldn't multiply 3 by -1 for some reason